what is the expansion of ln(x+1)

Grade:11

2 Answers

ruchi yadav
10 years ago
ln(x+1) = x-(1/2x2) +(1/3x3)-(1/4x4).....for -1<x<1

Thank You
Ruchi
Askiitians Faculty
sivakrishna
10 years ago
ln(x+1)=1/(x+1)